OpenStudy (anonymous):

If f(x) = 9 – 2|x + 2|, find f(–7). @gigi.antonia @pgpilot326

OpenStudy (anonymous):

I just dont understand the f and x part and how I solve with them.

OpenStudy (anonymous):

f is your function (f of x) so whenever you have f of something, you replace all of the exes with the number it tells you to. f(-7)=9-2l-7+2l

OpenStudy (anonymous):

so 35?

OpenStudy (anonymous):

Im so confused!

OpenStudy (anonymous):

think of the x as being in parentheses... in your case, \(f(x)=9-2|x+2|\) think of it like this... \[f(x)=9-2|(x)+2|\] so whatever is in the parentheses on the left is what should be in the parenetheses on the right. in your case... \[f(-7)=9-2|(-7)+2|\] now just use the order of operations to resolve and get the value of f(-7).

OpenStudy (anonymous):

9-2 is 7 then -7+ 2 is -5 but the absolute value is 5 but then what do I do?

OpenStudy (anonymous):

work it.. f(-7)=9-2(5)

OpenStudy (anonymous):

remember the order of operations

OpenStudy (anonymous):

Think of the absolute value sign as parentheses

OpenStudy (anonymous):

so -1?

OpenStudy (anonymous):

yep... 9-10= -1

OpenStudy (anonymous):

oh wow thanks I think I'm grasping the concept! If I need help in a few minutes will uyou be willing to help?

OpenStudy (anonymous):

yep should be around

OpenStudy (anonymous):

Same here

OpenStudy (anonymous):

ok thanks!!!!